4. A liquid of specific heat capacity of 3000 JKg-1K
-1
rise from 15 oC to 65 oC in one minute
when an electric heater is used. If the heater generates 63000 J per minutes, calculate
the mass of the water.
1
Expert's answer
2020-07-08T08:23:57-0400

The heat "Q" generated in one minute is:

"Q = 63000(J\/min)\u00b71(min) = 63000 (J)".

Assuming that no heat was lost , the heat generated equals the heat absorbed by the liquid.

The relation between the heat "Q" absorbed by the liquid and the temperature change "T_2-T_1" is the following:

"Q = cm(T_2-T_1)" ,

where "c" is the specific heat capacity of the liquid and "m" is its mass. The temperature change is the difference temperature and has the same numerical value in °C and in K, therefore "T_2-T_1 = 65-15 = 50" K.

Making use of this relation, we can deduce the mass of the liquid:

"m = \frac{Q}{c(T_2-T_1)} = \frac{63000 (J)}{3000(Jkg^{-1}K^{-1})·(50) (K)} = 0.42" kg.

Answer: the mass of the water is 0.42 kg.